Find the taylor series for sin x at a=Ï€
1
Expert's answer
2020-05-20T19:25:27-0400

"f(x)=\sum\limits_{i=1}^{\infty}f^{(i)}(a)/i!*(x-a)^i"

"sin(\pi)=\sum\limits_{i=1}^{\infty}f^{(2*i-1)}(\pi)/(2*i-1)!*(x-\pi)^{2*i-1}"

The Taylor Series of sin(x) with center π : -(x-π)+1/6*(x-π)^3-1/120*(x-π)^5+1/5040*(x-π)^7+...